OpenStudy (anonymous):

what if you scored an 0.80 on a test what fraction of the test in simplest form did you answer correctley

jimthompson5910 (jim_thompson5910):

what decimal place is the 8 in?

jimthompson5910 (jim_thompson5910):

if you don't know, you can say so

OpenStudy (anonymous):

tenths

jimthompson5910 (jim_thompson5910):

so we can say 0.8 is the same as 8 tenths = 8/10 now reduce 8/10 to get ______ (fill in the blank)

OpenStudy (anonymous):

what

jimthompson5910 (jim_thompson5910):

are you able to reduce 8/10 ?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

what are some common factors between 8 and 10?

OpenStudy (jacksonjrb):

In other words, what is the greatest common multiple between the two?

OpenStudy (anonymous):

40

OpenStudy (jacksonjrb):

What divides into both equally?

jimthompson5910 (jim_thompson5910):

you're thinking the LCM

jimthompson5910 (jim_thompson5910):

factors of 8: 1, 2, 4, 8 factors of 10: 1, 2, 5, 10 are you able to spot the GCF?

OpenStudy (anonymous):

ohhh um 2

OpenStudy (jacksonjrb):

So divide 8 and 10 by 2.

jimthompson5910 (jim_thompson5910):

yes so you divide both parts of 8/10 by 2 8/2 = 4 10/2 = 5 this means 8/10 = 4/5